# Computing Certificates of Strictly Positive Polynomials in Archimedean Quadratic Modules

### v1

@@ 3.2 Generalization of Averkov’s Lemma
when $N>\frac{\log(\Lambda)+2k\log(\gamma+\epsilon)-\log(\gamma)}{2\log(2)}+k$. Note that the right-hand side of the above inequality is greater than 00, which implies that $f-h>0$. ∎
@@
Consider the polynomial $f=9-x^{3}-y^{3}+3xy-5x^{2}$ and an Archimedean polynomial $g=1-x^{2}-y^{2}$ in $\mathbb{R}[x,y]$. One can check that $\{(x,y):f(x,y)\leq 0\}\cap\{(x,y):g(x,y)\geq 0\}=\emptyset$, and thus $f$ is strictly positive over $\operatorname{S}(g)$.
Let $B$ in Theorem 3.3 be $\mathbb{R}^{2}$ itself. Then assign the parameters $\gamma=\frac{2}{3}$, $\epsilon=\frac{1}{6}$, and $N=2$ to make $f-\tilde{g}(\frac{\tilde{g}-\gamma}{\gamma+\epsilon})^{N}$ strictly positive on $\mathbb{R}^{2}$: one can check that $\{(x,y):f(x,y)-\tilde{g}(\frac{\tilde{g}-\gamma}{\gamma+\epsilon})^{N}(x,y)\leq 0\}=\emptyset$.
The proof of Theorem 3.3 is a constructive one, that is to say, one can explicitly construct the polynomial $h=(\sum_{i=1}^{n}x_{i}^{2})(\frac{\tilde{g}-\gamma}{\gamma+\epsilon})^{2N}g\in\operatorname{QM}(g)$ via construction of the intermediate polynomial $\tilde{g}$ and computation of involved parameters $\epsilon$, $\gamma$, and $N$. The values of $\epsilon$ and $\gamma$ are the same as in (2), but that of $N$ needs to satisfy two inequalities in the two cases and is dependent on other parameters like $M$, $K$, and $A$ in the proof, and thus its computation is a bit complex. But if we endow $f$ in Theorem 3.3 with an additional property to have a lower bound on $B$, then the proof of Theorem 3.3 will become much more simplified in the following two aspects: (i) The explicit construction of $K$ and $\tilde{g}$ becomes unnecessary. Instead, the polynomial $g$ directly replaces $\tilde{g}$. (ii) The analysis with the set $D$ can be omitted, so is case (2). This is because, in case (1), for any $\bar{\bm{x}}\in B$, when $g(\bar{\bm{x}})<-2\epsilon$, the finiteness of $L:=\min_{\bm{y}\in B}f(\bm{y})$ implies $f(\bar{\bm{x}})-h(\bar{\bm{x}})<L+C(N)$ which covers case (2).
We formalize these observations as the following Theorem 3.5 and omit its formal proof. In fact, the simplifications we discuss above due to the existence of the lower bound make the proof of Theorem 3.5 quite similar to that of Lemma 2.3: the polynomial $h$ is in the form $h=(\frac{g-\gamma}{\gamma+\epsilon})^{2N}g$ and the values of all the parameters are the same as in (2).
@@
Let $g\in\mathbb{R}[\bm{x}]$ be a polynomial with bounded $\operatorname{S}(g)$ and $f\in\mathbb{R}[\bm{x}]$ be strictly positive on $\operatorname{S}(g)$. Then, for any subset $B$ of $\mathbb{R}^{n}$, if $f$ has a lower bound on $B$, there exists a polynomial $h\in\operatorname{QM}(g)$ such that $f-h>0$ over B.
Comparing Theorems 3.3 and 3.5, one can see that the former is more general, while the latter, with an additional assumption, is easier to implement. In our method for solving the certificate problem, we mainly use Theorem 3.5 and below we formulate it into Algorithm 2. In fact, our experiments show that using Theorem 3.5 by pre-construction of another polynomial with a lower bound usually leads to lower degrees of the computed certificates than using Theorems 3.3 directly.
Input: A polynomial $g\in\mathbb{R}[\bm{x}]$ with bounded $\operatorname{S}(g)$, a polynomial $f\in\mathbb{R}[\bm{x}]$ which has a lower bound on $B$ and is strictly positive over $\operatorname{S}(g)$, and a set $B\subseteq\mathbb{R}^{n}$
Output: A sum-of-squares multiplier $\sigma$ in $\mathbb{R}[\bm{x}]$ such that $f-\sigma g>0$ over $B$
@@ 8       Choose $\gamma$, $\epsilon$, and $N$ as in (2);
4 Extending Lasserre’s method to a monogenic Archimedean quadratic module
Lasserre gave a method in [6] to approximate a nonnegative polynomial by using sums of squares, and it shows that every real nonnegative polynomial $f$ can be approximated as closely as desired by a sequence of polynomials that are sums of squares. This method is extended to work on a monongenic Archimedean quadratic module. Lasserre’s result is reviewed below first; this is followed by the proposed extension.
Theorem 4.1 ([6, Theorem 4.1]).
Let $f\in\mathbb{R}[\bm{x}]$ be nonnegative with global minimum $f^{*}$, that is, $f(\bm{x})\geq f^{*}\geq 0$ for any $\bm{x}\in\mathbb{R}^{n}$. Then for every $\epsilon>0$, there exists some $r(\epsilon,f)\in\mathbb{N}$ such that,
@@ 4 Extending Lasserre’s method to a monogenic Archimedean quadratic module
is also a sum of squares. The parameter $r$ mentioned can be determined by solving a classical semi-definite programming problem. For the details in this method, one may refer to [6].
Next we extend this method to a monogenic quadratic module $\operatorname{QM}(g)$ with a bounded $\operatorname{S}(g)$. For a positive polynomial $f$, our objective is to identify an element $h\in\operatorname{QM}(g)$ with $\operatorname{S}(g)$ bounded such that $f-h$ is a sums of squares over $\mathbb{R}^{n}$.
@@
Let $f\in\mathbb{R}[\bm{x}]$ be a positive polynomial with a global minimum $f^{*}>0$ and $g\in\mathbb{R}[\bm{x}]$ with $\operatorname{S}(g)$ bounded. Then there exist $\epsilon>0$ and $r\in\mathbb{N}$ such that
$f_{\epsilon}=f-\epsilon g\sum_{k=0}^{r}\sum_{j=1}^{n}\frac{x_{j}^{2k}}{k!}$ (3)
@@
Denote the sum of squares $\sum_{k=0}^{r}\sum_{j=1}^{n}\frac{x_{j}^{2k}}{k!}$ by $p_{r}(\bm{x})$. In the proof of Proposition 3.3, we show that there exists $K>0$ such that $\operatorname{S}(g+K)$ is bounded, i.e. there exists $R>0$ such that $\operatorname{S}(g+K)\subseteq\{\bar{\bm{x}}\leavevmode\nobreak\ |\leavevmode\nobreak\ \|\bar{\bm{x}}\|^{2}\leq R\}$. Then, for any $\bar{\bm{x}}$ such that $\|\bar{\bm{x}}\|^{2}\leq R$, we have $p_{r}(\bar{\bm{x}})\leq\sum_{k=0}^{r}\sum_{j=1}^{n}\frac{R^{k}}{k!}\leq ne^{R}$ for any $r>0$.
Consider the polynomial
$f_{\epsilon}:=f-\epsilon gp_{r}=(f-\epsilon(g+K)p_{r})+\epsilon Kp_{r}.$
Let $M=\max_{\bm{x}}(g+K)$, which is finite, and assign $\epsilon=\frac{f^{*}}{2Mne^{R}}$. Then we have $h:=f-\epsilon(g+K)p_{r}\geq f-\frac{f^{*}}{2}>0$. Applying Theorem 4.1 to $h$ with $\epsilon K$, we know that there exists some $r(\epsilon K,h)\in\mathbb{N}$ such that $h+\epsilon Kp_{r(\epsilon K,h)}=f_{\epsilon}$ is a sum of squares. ∎
This constructive proof immediately translates to an algorithmic procedure we formulate as Algorithm 3. Instead of computing $r$ by solving a SDP problem in the algorithm, one can simply increase the value of $r$ iteratively and check whether the resulting $f_{\epsilon}=f-\epsilon gp_{r}$ becomes a sum of squares until it does.
Input: A polynomial $g\in\mathbb{R}[\bm{x}]$ with bounded $\operatorname{S}(g)$, a polynomial $f\in\mathbb{R}[\bm{x}]$ such that $f>0$ over $\mathbb{R}^{n}$
Output: A sum-of-squares multiplier $\sigma$ in $\mathbb{R}[\bm{x}]$ such that $f-\sigma g$ is a sum of squares
@@
5       Choose $K$ and $R$ in proof of Theorem 4.2;
6       $f^{*}=\min_{\bm{x}}f$; $M=\max_{\bm{x}}(g+K)$;
@@ 8       Compute $r$ by solving a SDP problem ;
9       return $\epsilon\sum_{k=0}^{r}\sum_{j=1}^{n}\frac{x_{j}^{2k}}{k!}$;
10
@@
Consider the polynomials $g=1-x^{4}y^{2}-x^{2}y^{4}+x^{2}y^{2}-y^{6}-x^{6}$ and $f=2+x^{4}y^{2}+x^{2}y^{4}-3x^{2}y^{2}$. The latter is the Motzkin’s polynomial plus 1, which is strictly positive over $\mathbb{R}^{2}$. Using the technique in [3, Lemma 1], one can show that $f$ is not a sum of squares. Choosing $K=1$, the minimum value of $f$ is $1$ and the maximum value of $g+K$ is $\frac{217}{108}$. Then the value of $\epsilon$ is $\frac{27}{217e^{3}}$. We start with $r=0$: now $p_{r}=n$ and $h=f-\epsilon(g+K)n$, and in fact $f_{\epsilon}=h+\epsilon Kn$ is verified to be a sum of squares.
5 Computing certificates using extended Averkov construction and Lasserre’s method
In this section, the extended Averkov construction and Lasserre’s method are integrated to compute certificates in the general case. There are two issues which need to be addressed; this is discussed in the next subsection. An algorithm to compute certificates is given in the following subsection.
@@ 5 Computing certificates using extended Averkov construction and Lasserre’s method
Two issues need to be addressed to apply the extended Averkov’s meth-od. Per Theorem 3.5, a polynomial $g\in\operatorname{QM}(G)$ with a bounded $\operatorname{S}(g)$ must be constructed; further, its global lower bound needs to be computed.
The univariate and multivariate cases are handled separately for computing the above $g$. If there exists a generator $g_{i}\leavevmode\nobreak\ (1\leq i\leq s)$ of $\operatorname{QM}(G)$ such that $\operatorname{S}(g_{i})$ is bounded, then use that $g_{i}$ as $g$. Otherwise, assign $g$ to be the Archimedean polynomial $g:=N-\sum x_{i}^{2}$ with a proper choice of $N$ such that $f>0$ on $\operatorname{S}(g)$. In case every $\operatorname{S}(g_{1}),\ldots,\operatorname{S}(g_{s})$ unbounded, in the univariate case, there must exist two generators $g_{i}$ and $g_{j}\leavevmode\nobreak\ (i\neq j)$ such that the leading coefficients $\operatorname{lc}(g_{i})$ and $\operatorname{lc}(g_{j})$ have opposite signs and the degrees of $g_{i}$ and $g_{j}$ are odd. Let
$\left\{\begin{array}[]{ll}c_{i}:=\frac{1}{|\operatorname{lc}(g_{i})|},c_{j}:=\frac{(x-a)^{\deg(g_{i})-\deg(g_{j})}}{|\operatorname{lc}(g_{j})|},&\mbox{if}\deg(g_{i})>\deg(g_{j}),\\ c_{i}:=\frac{x^{2}}{|\operatorname{lc}(g_{i})|},c_{j}:=\frac{(x-a)^{2}}{|\operatorname{lc}(g_{j})|},&\mbox{if}\deg(g_{i})=\deg(g_{j}),\end{array}\right.$ (4)
where the constant $a$ is selected such that $\operatorname{lc}(g)<0$ and $g$ is of an even degree. Then the polynomial $g:=c_{i}g_{i}+c_{j}g_{j}$ is obtained by canceling the leading odd terms of $g_{i}$ and $g_{j}$. Clearly, we have $\operatorname{S}(G)\subseteq\operatorname{S}(g)$ and $g\in\operatorname{QM}(G)$.
The next step is to to construct a new polynomial $\tilde{f}:=f-\sigma g$ such that $\tilde{f}>0$ on $\operatorname{S}(g)$, where $\sigma$ is the certificate returned by $\operatorname{\sf Averkov}(\{g\},f,\operatorname{S}(g))$. At this point $\tilde{f}$ and $g$ as constructed above, satisfy the conditions of Theorem 3.3, but using Theorem 3.5 $\tilde{f}$ still misses a critical property of possessing a lower bound, in particular a global lower bound if $B$ in Theorem 3.5 is $\mathbb{R}^{n}$. The following proposition addresses this problem.
@@
Let $f,g\in\mathbb{R}[\bm{x}]$ be two polynomials with $\operatorname{S}(g)$ bounded and $f>0$ on $\operatorname{S}(g)$. Then there exists a sum of squares $\delta$ such that $f-\delta g$ is strictly positive on $\operatorname{S}(g)$ and has a lower bound over $\mathbb{R}^{n}$.

### v2

@@ 3.2 Generalization of Averkov’s Lemma
when $N>\frac{\log(\Lambda)+2k\log(\gamma+\epsilon)-\log(\gamma)}{2\log(2)}+k$. Note that the right-hand side of the above inequality is greater than 0, which implies that fh>0f-h>0. ∎
Remark 3.6.
Since the set $B$ in Lemma LABEL:lemma:c is compact, it is proved that the polynomial obtained fhf-h is not only strictly positive on $B$ but it also has a strictly positive minimum on $B$. In fact, the polynomial fhf-h in Theorem LABEL:lemma:d can also be proven to have a strictly positive minimum on $B$, even though $B$ here is not necessarily compact: in part (2) of the proof, let $\mu=-\Lambda+\gamma\left(\frac{1}{\gamma+\epsilon}\right)^{2k}\left(\frac{2\gamma+2\epsilon}{\gamma+\epsilon}\right)^{2N-2k}>0$. Then we have $f-h>\mu\tilde{g}^{2k}\to+\infty$ when $\|\bm{x}\|\to\infty$. This property of fhf-h will be used later in combination with Theorem LABEL:avoid_archimedean_poly in our algorithm to compute the certificates.
If the polynomial $g$ in the theorem statement already satisfies the inequality (5), we can directly set $\tilde{g}:=g$. Specifically, all Archimedean polynomials satisfy this inequality.
[…]
@@
Consider the polynomial $f=9-x^{3}-y^{3}+3xy-5x^{2}$ and an Archimedean polynomial $g=1-x^{2}-y^{2}$ in $\mathbb{Q}[x,y]$. One can check that $\{(x,y):f(x,y)\leq 0\}\cap\{(x,y):g(x,y)\geq 0\}=\emptyset$, and thus $f$ is strictly positive over $\operatorname{S}(g)$.
Let $B$ in ?THM? LABEL:lemma:d be $\mathbb{R}^{2}$ itself and $\tilde{g}=g$. Then, we assign the parameters $\gamma=\frac{2}{3}$, $\epsilon=\frac{1}{6}$, and $N=2$ to make $f-(\frac{\tilde{g}-\gamma}{\gamma+\epsilon})^{N}\cdot\tilde{g}$ strictly positive on $\mathbb{R}^{2}$: one can check that $\{(x,y):f(x,y)-(\frac{\tilde{g}(x,y)-\gamma}{\gamma+\epsilon})^{N}\cdot\tilde{g}(x,y)\leq 0\}=\emptyset$.
The proof of ?THM? LABEL:lemma:d is a constructive one, that is to say, one can explicitly construct the polynomial $h=(\frac{\tilde{g}-\gamma}{\gamma+\epsilon})^{2N}\cdot\tilde{g}\in\operatorname{QM}(g)$ via construction of the intermediate polynomial $\tilde{g}$ and the computation of the involved parameters $\epsilon$, $\gamma$, and $N$. The values of $\epsilon$ and $\gamma$ are the same as in (2), but $N$ needs to satisfy two inequalities in the two cases and is dependent on other parameters like $M$, and $A$ in the proof, and thus its computation is a bit complex. But if we endow $f$ in Theorem LABEL:lemma:d with an additional property to have a lower bound on $B$, then the proof of Theorem LABEL:lemma:d will become much more simplified in the following two aspects: (i) The explicit construction of $\tilde{g}$ becomes unnecessary. Instead, the polynomial $g$ directly replaces $\tilde{g}$. (ii) The analysis with the set $D$ can be omitted, so is case (2). This is because, in case (1), for any $\bar{\bm{x}}\in B$, when $g(\bar{\bm{x}})<-2\epsilon$, the finiteness of $L:=\min_{\bar{\bm{x}}\in B}f(\bar{\bm{x}})$ implies $f(\bar{\bm{x}})-h(\bar{\bm{x}})<L+C(N)$ which covers case (2).
We formalize these observations as the following ?THM? LABEL:cor:lowerbound and omit its formal proof. In fact, the simplifications we discuss above due to the existence of the lower bound make the proof of ?THM? LABEL:cor:lowerbound quite similar to that of ?THM? LABEL:lemma:c: the polynomial $h$ is in the form $h=(\frac{g-\gamma}{\gamma+\epsilon})^{2N}\cdot g$ and the values of all the parameters are the same as in (2).
@@
Let $g\in\mathbb{R}[X]$ be a polynomial with bounded $\operatorname{S}(g)$ and $f\in\mathbb{R}[X]$ be strictly positive on $\operatorname{S}(g)$. Then, for any subset $B$ of $\mathbb{R}^{n}$, if $f$ has a lower bound on $B$, there exists a polynomial $h\in\operatorname{QM}(g)$ such that fh>0f-h>0 over B.
Comparing ?THM? LABEL:lemma:d and ?THM? LABEL:cor:lowerbound, one can see that the former is more general, while the latter, with an additional assumption, is easier to implement. In our method for solving the certificate problem, we mainly use ?THM? LABEL:cor:lowerbound and below we formulate it into Algorithm 2. In fact, our experiments show that using ?THM? LABEL:cor:lowerbound by pre-construction of another polynomial with a lower bound usually leads to lower degrees of the computed certificates than using ?THM? LABEL:lemma:d directly.
Input: A polynomial $g\in\mathbb{K}[X]$ with bounded $\operatorname{S}(g)$, a polynomial $f\in\mathbb{K}[X]$ which has a lower bound on $B$ and is strictly positive over $\operatorname{S}(g)$, and a set $B\subseteq\mathbb{R}^{n}$
Output: A sum-of-squares multiplier $\sigma$ in $\mathbb{K}[X]$ such that $f-\sigma\cdot g>0$ over $B$
@@ 6   Choose $\gamma$, $\epsilon$, and $N$ as in (2);
4 Extending Lasserre’s method to a monogenic Archimedean qua-dratic module
Lasserre (2006) gave a method to approximate a non-negative polynomial by using sums of squares, and it shows that every real non-negative polynomial $f$ can be approximated as closely as desired by a sequence of polynomials that are sums of squares. This method is extended to work on a monogenic Archimedean quadratic module. Lasserre’s result is recalled below first; this is followed by the proposed extension.
Theorem 4.1 (Lasserre (2006, Theorem 4.1)).
Let $f\in\mathbb{R}[X]$ be non-negative with global infimum $f^{*}=\inf_{\bar{\bm{x}}\in\mathbb{R}^{n}}f(\bar{\bm{x}})$, that is, $f(\bar{\bm{x}})\geq f^{*}\geq 0$ for any $\bar{\bm{x}}\in\mathbb{R}^{n}$. Then for every $\epsilon>0$, there exists some $r(\epsilon,f)\in\mathbb{N}$ such that,
@@ 4 Extending Lasserre’s method to a monogenic Archimedean qua-dratic module
is also a sum of squares. The parameter $r$ mentioned can be determined by solving a classical semi-definite programming problem. For details of this method, we refer to Lasserre (2006).
Next, we extend this method to a monogenic quadratic module $\operatorname{QM}(g)$ with a bounded $\operatorname{S}(g)$. For a positive polynomial $f$, our objective is to identify an element $h\in\operatorname{QM}(g)$ with $\operatorname{S}(g)$ bounded such that fhf-h is a sum of squares over $\mathbb{R}^{n}$.
@@
Let $f\in\mathbb{R}[X]$ be a positive polynomial with a global minimum $f^{*}=\inf_{\bar{\bm{x}}\in\mathbb{R}^{n}}f(x)>0$ and $g\in\mathbb{R}[X]$ with $\operatorname{S}(g)$ bounded. Then there exist $\epsilon>0$ and $r\in\mathbb{N}$ such that
$f_{\epsilon}=f-\epsilon\sum_{k=0}^{r}\sum_{j=1}^{n}\frac{x_{j}^{2k}}{k!}\cdot g$ (6)
@@
Denote the sum of squares $\sum_{k=0}^{r}\sum_{j=1}^{n}\frac{x_{j}^{2k}}{k!}$ by $p_{r}(X)$. Consider $\tilde{g}$ in the proof of ?THM? LABEL:lemma:d, then we show there that for any $K>0$, $\operatorname{S}(\tilde{g}+K)$ is bounded, i.e. there exists $R>0$ such that $\operatorname{S}(\tilde{g}+K)\subseteq\{\bar{\bm{x}}\penalty 10000\ |\penalty 10000\ \|\bar{\bm{x}}\|^{2}\leq R\}$. Then, for any $\bar{\bm{x}}$ such that $\|\bar{\bm{x}}\|^{2}\leq R$, we have $p_{r}(\bar{\bm{x}})\leq\sum_{k=0}^{r}\sum_{j=1}^{n}\frac{R^{k}}{k!}\leq ne^{R}$ for any $r>0$.
For the sake of simplicity, let $K=1$. Consider the polynomial
$f_{\epsilon}:=f-\epsilon p_{r}\cdot\tilde{g}=(f-\epsilon(\tilde{g}+1)p_{r})+\epsilon p_{r}.$
Let $M=\max_{\bar{\bm{x}}}(\tilde{g})+1$, which is finite, and assign $\epsilon=\frac{f^{*}}{2Mne^{R}}$. Then we have $h:=f-\epsilon(\tilde{g}+1)p_{r}\geq f-\frac{f^{*}}{2}>0$. Applying ?THM? LABEL:lasserre to $h$ with $\epsilon$, we know that there exists some $r(\epsilon,h)\in\mathbb{N}$ such that $h+\epsilon p_{r(\epsilon,h)}=f_{\epsilon}$ is a sum of squares. ∎
This constructive proof immediately translates to an algorithmic procedure we formulate as Algorithm 3. Instead of computing $r$ by solving a SDP problem in the algorithm, one can simply increase the value of $r$ iteratively and check whether the resulting $f_{\epsilon}=f-\epsilon p_{r}\cdot\tilde{g}$ becomes a sum of squares.
Input: A polynomial $g\in\mathbb{K}[X]$ with bounded $\operatorname{S}(g)$, a polynomial $f\in\mathbb{K}[X]$ such that $f>0$ over $\mathbb{R}^{n}$
Output: A sum-of-squares multiplier $\sigma$ in $\mathbb{K}[X]$ such that $f-\sigma\cdot g$ is a sum of squares
@@
5   Choose $R$ in proof of ?THM? LABEL:avoid_archimedean_poly;
6 if $\operatorname{S}(g+1)$ is not bounded then
7    $c=\|X\|^{2\lfloor\frac{(d-1)^{n}}{2}\rfloor+2}$;
8 else
9    $c=1$;
10 $f^{*}=\min_{\bar{\bm{x}}}f$; $M=\max_{\bar{\bm{x}}}(c\cdot g)+1$;
@@ 12   Compute $r$ by solving a SDP problem ;
13 return $c\cdot\epsilon\sum_{k=0}^{r}\sum_{j=1}^{n}\frac{x_{j}^{2k}}{k!}$;
14
@@
Consider the polynomials $g=1-x^{4}y^{2}-x^{2}y^{4}+x^{2}y^{2}-y^{6}-x^{6}$ and $f=2+x^{4}y^{2}+x^{2}y^{4}-3x^{2}y^{2}$. The latter is the Motzkin’s polynomial plus 1, which is strictly positive over $\mathbb{R}^{2}$. Using the technique in Berg (1987, Lemma 1), one can show that $f$ is not a sum of squares.The minimum value of $f$ is $1$ and the maximum value of $g+1$ is $\frac{217}{108}$. Then the value of $\epsilon$ is $\frac{27}{217e^{3}}$. We start with $r=0$: now $p_{r}=n$ and $h=f-\epsilon(g+1)n$, and in fact $f_{\epsilon}=h+\epsilon n$ is verified to be a sum of squares.
5 A constructive proof of Putinar’s criterion
With the necessary extensions now in place to address the certificate problem, we apply the constructive approach here to prove ?THM? LABEL:thm:archimedean, by computing the certificate of Archimedean polynomial in terms of $\operatorname{QM}(g)$ with bounded $g$.
Proof of ?THM? LABEL:thm:archimedean.
[…]
$f=(h+\epsilon p_{r})\cdot g+\tilde{f}_{\epsilon}.$ (7)
[…]
6 Computing certificates using the extended Averkov construction and Lasserre’s method
In this section, the extended Averkov construction and Lasserre’s method are integrated to compute certificates in the general case. There are two issues that need to be addressed; this is discussed in the next subsection. An algorithm to compute certificates is given in the subsection to follow.
@@ 6 Computing certificates using the extended Averkov construction and Lasserre’s method
Two issues need to be addressed to apply the extended Averkov’s method. Per ?THM? LABEL:cor:lowerbound, a polynomial $g\in\operatorname{QM}(G)$ with a bounded $\operatorname{S}(g)$ must be constructed; further, its global lower bound needs to be computed.
[…]
$\left\{\begin{array}[]{ll}c_{i}:=\frac{1}{|\operatorname{lc}(g_{i})|},c_{j}:=\frac{(x-a)^{\deg(g_{i})-\deg(g_{j})}}{|\operatorname{lc}(g_{j})|},&\mbox{if\penalty 10000\ }\deg(g_{i})>\deg(g_{j}),\\ c_{i}:=\frac{x^{2}}{|\operatorname{lc}(g_{i})|},c_{j}:=\frac{(x-a)^{2}}{|\operatorname{lc}(g_{j})|},&\mbox{if\penalty 10000\ }\deg(g_{i})=\deg(g_{j}),\end{array}\right$. (8)
The constant $a$ is selected such that $\operatorname{lc}(g)<0$. One can see that the choice of $c_{i},c_{j}$ forces the cancellation of the leading odd terms of $g_{i}$ and $g_{j}$, hence $\deg(g)$ is even. Clearly, we have $\operatorname{S}(G)\subseteq\operatorname{S}(g)$ and $g\in\operatorname{QM}(G)$.
The next step is to construct a new polynomial $\tilde{f}:=f-\sigma\cdot g$ such that $\tilde{f}>0$ on $\operatorname{S}(g)$, where $\sigma$ is the certificate returned by $\operatorname{\sf Averkov}(\{g\},f,\operatorname{S}(g))$. At this step, $\tilde{f}$ and $g$ as constructed above satisfy the conditions of ?THM? LABEL:lemma:d; however, $\tilde{f}$ still misses a critical property of possessing a lower bound, in particular a global lower bound if $B$ in ?THM? LABEL:cor:lowerbound is $\mathbb{R}^{n}$. The following proposition addresses this problem.
@@
Let $f,g\in\mathbb{R}[X]$ be two polynomials with $\operatorname{S}(g)$ bounded and $f>0$ on $\operatorname{S}(g)$. Then there exists a sum of squares $\delta$ such that $f-\delta\cdot g$ is strictly positive on $\operatorname{S}(g)$ and has a lower bound over $\mathbb{R}^{n}$.